a sofa is 12 yrs. old as teble is 36 yrs. old . in how many years will the table be twice as old as sofa

Answer by suruman(21)

A sofa is currently 12 years old. A table is currently 36 years old.
After x years from now, table will be twice as old as sofa.
36 + x = 2* (12 + x)
36 + x = 24 + 2x
12 = x
Therefore, after 12 years from now, Table will be twice as old as Sofa.